What is the energy of a photon with a frequency of 7.17 X 10^14 s^-1? (h=6.626 x 10^-34 J X S)
1
Expert's answer
2020-04-28T12:20:38-0400
"E=h\nu"

"E = 6.626*10^{-34}*7.17*10^{14} = 47.5*10^{-20}J"